An Individual Retirement Account​ (IRA) has ​$9 comma 000 in​ it, and the owner decides not to add any more money to the account other than interest earned at 8 ​% compounded daily. How much will be in the account 29 years from now when the owner reaches retirement​ age? Round to the nearest cent. Use a​ 365-day year.
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=9000*(1+0.08/365)^(365*29)

=9000*10.17

=$91,530.